### v1

@@ 1 Introduction
Quasi-Monte Carlo (QMC) methods replace random sampling with carefully constructed deterministic point sets for numerical integration over the unit cube; see, e.g., [4, 5, 6, 13, 14, 19]. While classical QMC theory focuses on discrepancy, which measures deviation from perfect equidistribution, it does not directly control local spacing between points. However, many applications require a more geometric notion of uniformity. In particular, tasks such as scattered data approximation [23], Gaussian process regression [22], kernel interpolation [18], and the design of computer experiments [7, 15, 17] benefit from point sets that ensure both well-controlled spacing and coverage. These properties are summarized by the term quasi-uniformity.
@@
Intuitively, $h_{p}(Q)$ is the smallest radius such that closed $\ell^{p}$ balls around each point cover $0,1]^{d}$, whereas $q_{p}(Q)$ is the largest radius such that the corresponding open balls around the points do not overlap. As noted in [14, Section 6], [21], and [16], this geometric interpretation implies
@@
Given that QMC point sets are natural candidates for generating uniform point sets, it is important to ask whether they are quasi-uniform. As noted in [24], however, the quasi-uniformity of classical QMC constructions had long been unresolved. This changed with Goda’s influential result [9], which shows that the two-dimensional Sobol’ sequence is not quasi-uniform.
The Sobol’ sequence [20] is widely used in practice due to its efficient digital construction, extensibility to arbitrary sample sizes, and effectiveness in high-dimensional integration. It is available in standard software packages, such as Python’s QMCPy [1] and MATLAB’s Statistics and Machine Learning Toolbox. Sobol’ and Shukhman [21] conjectured that the separation radius of the first $N$ points of the $d$-dimensional Sobol’ sequence behaves like $N^{-1/d}$, which would imply quasi-uniformity. For $d=2$, Goda [9] showed that this conjecture fails for the sparse subsequence of the form $N=2^{2^{w}-1}$, and the case $N=2^{2^{w}}$ was subsequently analyzed in [3]. Nevertheless, a complete description of the separation radius for general $N$ remains open, and addressing this gap is the focus of the present work.
@@
  • We provide an exact formula for the $\ell^{\infty}$-separation radius of the dyadic prefixes $Q_{2^{m}}$.
@@
Otherwise, decomposing $m$ as $m=2^{v}+2^{w}+c$ with integers $v>w\geq 0$ and $0\leq c<2^{w}$, then
@@
$q_{\infty}(Q_{2^{m}})$ $\leq 2^{-3m/4-3/2}\quad\text{for }m\neq 1,5,$ (3)
@@
Related work. In recent years, the quasi-uniformity of QMC point sets has been the subject of intensive investigation. There are two major classes of QMC constructions: digital nets and sequences [6, 14], and lattice point sets (including their infinite analogue, Kronecker sequences) [4, 19].
For lattice point sets, quasi-uniformity has been actively studied. In one dimension, the separation radius of Kronecker sequences is completely characterized in [8]. More general results in higher dimensions are provided in [2], including bounded mesh ratios for two-dimensional Fibonacci lattices, existence results for $d$-dimensional lattice rules, and explicit constructions for $d$-dimensional Kronecker sequences.
Turning to digital nets and sequences, the one-dimensional van der Corput sequence in base $b$ is known to be quasi-uniform since its first $b^{m}$ points are $\{i/b^{m}\mid 0\leq i<b^{m}\}$. In higher dimensions, the covering radius, often referred to as dispersion, has been extensively studied [14, Chapter 6]. In particular, for $(t,d)$-sequences in base bb—well-known examples include the Sobol’, Faure, and Niederreiter sequences—, the covering radius is known to attain the optimal order $\Theta(N^{-1/d})$ for any dimension $d$. Thus, the problem of establishing quasi-uniformity reduces to verifying whether the separation radius also scales as $N^{-1/d}$.
For $d=2$, as stated, the Sobol’ sequence is not quasi-uniform [9]. The separation radius of several two-dimensional digital nets was studied in [10, 11]. Numerical experiments therein suggest that the Larcher–Pillichshammer nets [12] are quasi-uniform. This was theoretically proved by Dick, Goda, and Suzuki [3], who introduced an algebraic criterion for well-separated digital nets. To our knowledge, this remains the only explicit construction of low-discrepancy and quasi-uniform digital nets for $d\geq 2$. The paper [3] also shows the non-optimality of the separation radius for some two-dimensional digital nets and Fibonacci polynomial lattices, as well as for $b$-dimensional Faure sequences in prime base $b$.
From a different perspective, Pronzato and Zhigljavsky [16] constructed quasi-uniform infinite sequences via a greedy packing algorithm, ensuring a mesh ratio of at most $2$ for the first $N$ points, $N\geq 2$. However, these sequences do not necessarily maintain low discrepancy in dimensions $d\geq 2$.
Organization. Preliminaries and notation are collected in Section 2. Section 3 provides the necessary lemmas, the proof of Theorem 1.1, and the derivation of Corollaries 1.2 and 1.3.
@@ 2 Preliminaries
For a vector $\boldsymbol{z}=(z_{1},\dots,z_{m})^{\top}\in\mathbb{F}_{2}^{m}$, we denote by $\boldsymbol{z}[i]\in\{0,1\}$ its $i$th component $z_{i}$, and for $i<j$ we define the slice $\boldsymbol{z}[i\!:\!j]=(z_{i},\dots,z_{j})^{\top}$. Finally, we set
$\phi(\boldsymbol{z})=\boldsymbol{z}[1]2^{-1}+\cdots+\boldsymbol{z}[m]2^{-m}.$
@@
Two-dimensional Sobol’ sequence. The two-dimensional Sobol’ sequence is defined as follows. For the definition in general dimension $d$, we refer the reader to [6, Chapter 8].
@@
Let $n\in\mathbb{N}_{0}$ and choose $m\in\mathbb{N}$ such that $n<2^{m}$. The two-dimensional Sobol’ sequence $(\boldsymbol{x}_{n})_{n\in\mathbb{N}}$ is a sequence of points in $0,1)^{2}$, where the $n$th point is given by
$\boldsymbol{x}_{n}:=(\phi(\vec{n}),\phi(P_{m}\vec{n})).$
This definition does not depend on the choice of $m$: increasing $m$ simply pads $\vec{n}$ with leading zeros, and since $P_{m}$ is upper triangular, $\phi(\vec{n})$ and $\phi(P_{m}\vec{n})$ remain unchanged.
This construction is generalized to the notions of digital nets and digital sequences. The uniformity of these point sets is usually measured by a quantity called the $t$-value (see [14, Chapter 4] or [6, Chapter 4]). It is known that the $t$-value of the two-dimensional Sobol’ sequence attains the best possible value, namely zero [6, Section 8.1]. This discussion can be formulated rigorously in the following proposition.
@@ 3.1 Lemmas
Using this fact, Items (vi) and (vii) also follow from Lucas’s theorem. ∎
@@ 3.2 Proof of the main theorem
  • The case $m=2^{v}-1$ is essentially treated in [9]. As shown in the proof of [9, Theorem 2.2], we have
    $\|\boldsymbol{x}_{1}-\boldsymbol{x}_{2^{m}-1}\|_{\infty}=2^{-m}.$
@@
  • The case $m=2^{v}$ is essentially treated in [3]. From the proof of [3, Theorem 4.2] with $L$ being the identity matrix, we deduce $\boldsymbol{x}_{2^{m}+1}=(1/2+1/2^{m},1/2+1/2^{m})$, which implies $\|\boldsymbol{x}_{1}-\boldsymbol{x}_{2^{m}+1}\|_{\infty}=2^{-m}$. Hence, in the same manner as the previous case, we obtain
@@ 3.2.1 Proof of the upper bound
Let $p=2^{V+W-1}+2^{W-1}$ and $q=2^{V+W}-2^{W}$. To prove $q_{\infty}(Q_{2^{m}})\leq 2^{-V-W}$, it suffices to show that $\|\boldsymbol{x}_{p}-\boldsymbol{x}_{q}\|_{\infty}=2^{-V-W+1}$.
First, we compute $\boldsymbol{x}_{p}=(x_{p,1},x_{p,2})$. Since $\vec{p}[i]=1$ if and only if $i=W$ or $i=V+W$, it follows from Lemma 3.1 (i) and (ii) that
$x_{p,1}$ $=\phi(\vec{p})=2^{-W}+2^{-V-W},$
$x_{p,2}$ $=\phi(P_{m}\vec{p})=\sum_{i=1}^{V+W}(P[i][W]\oplus P[i][V+W])2^{-i}=\sum_{i=V+1}^{V+W}2^{-i}=2^{-V}-2^{-V-W}.$
Next, we compute $\boldsymbol{x}_{q}=(x_{q,1},x_{q,2})$. Since $\vec{q}[i]=1$ for $W+1\leq i\leq V+W$, we have
$P_{m}\vec{q}$ $=P_{m}(\mathrm{concat}(\boldsymbol{1}_{W},\boldsymbol{0}_{m-W})\oplus\mathrm{concat}(\boldsymbol{1}_{V+W},\boldsymbol{0}_{m-V-W}))$
$=\mathrm{concat}(P_{W}\boldsymbol{1}_{W},\boldsymbol{0}_{m-W})\oplus\mathrm{concat}(P_{V+W}\boldsymbol{1}_{V+W},\boldsymbol{0}_{m-V-W}),$
@@ 3.2.2 Proof of the lower bound
Then, Corollary 3.3 (i) implies
@@
In this case, Corollary 3.3 (ii), together with the assumption $\phi(\vec{p})<\phi(\vec{q})$, implies
@@
In this case, using (8) and Lemma 3.1 (v), we have
$(P\vec{\Delta})[V+1\!:\!V+W-1]$ $=P[V+1\!:\!V+W-1][V+1\!:\!V+W-1]\cdot\vec{\Delta}[V+1\!:\!V+W-1]$
$\quad\oplus P[V+1\!:\!V+W-1][V+W\!:\!V+W]\cdot\vec{\Delta}[V+W]$
$\quad\oplus P[V+1\!:\!V+W-1][V+W+1\!:\!m]\cdot\vec{\Delta}[V+W+1\!:\!m]$
@@
$=\boldsymbol{1}\oplus P^{\prime}\cdot\vec{\Delta}[V+W+1\!:\!m],$
where $P^{\prime}:=P[V+1\!:\!V+W-1][V+W+1\!:\!m]$. Combined with (12), this gives
@@
Finally, (5) follows immediately from (4) together with the general bound $h_{\infty}(Q_{N})\geq 1/(2\sqrt{N})$ as given in [3, Remark 2.4]. ∎

### v2

@@ 1 Introduction
Quasi-Monte Carlo (QMC) methods replace random sampling with carefully constructed deterministic point sets for numerical integration over the unit cube; see, e.g., [DKP22, DKS13, DP10, LePi14, N92, SJ94]. While classical QMC theory focuses on discrepancy, which measures deviation from perfect equidistribution, it does not directly control local spacing between points. However, many applications require a more geometric notion of uniformity. In particular, tasks such as scattered data approximation [W05], Gaussian process regression [T20], kernel interpolation [Sch95], and the design of computer experiments [FLS06, PM12, SWN03] benefit from point sets that ensure both well-controlled spacing and coverage. These properties are summarized by the term quasi-uniformity.
@@
Intuitively, $h_{p}(Q)$ is the smallest radius such that closed $\ell^{p}$ balls around each point cover $0,1]^{d}$, whereas $q_{p}(Q)$ is the largest radius such that the corresponding open balls around the points do not overlap. As noted in [N92, Section 6], [SS07], and [PZ23], this geometric interpretation implies
@@
Given that QMC point sets are natural candidates for generating uniform point sets, it is important to ask whether they are quasi-uniform. This question is non-trivial, since quasi-uniformity does not necessarily imply low discrepancy, nor does low discrepancy imply quasi-uniformity for $d\geq 2$ [DGLPSxx, Section 1.3]. As noted in [WBG21], this issue remained open for a long time for classical QMC constructions. This situation changed with Goda’s influential result [G24a], which shows that the two-dimensional Sobol’ sequence is not quasi-uniform.
The Sobol’ sequence [So67] is widely used in practice due to its efficient digital construction, extensibility to arbitrary sample sizes, and effectiveness in high-dimensional integration. It is available in standard software packages, such as Python’s QMCPy [QMCpy] and MATLAB’s Statistics and Machine Learning Toolbox. Sobol’ and Shukhman [SS07] conjectured that the separation radius of the first $N$ points of the $d$-dimensional Sobol’ sequence behaves like $N^{-1/d}$, which would imply quasi-uniformity. For $d=2$, Goda [G24a] showed that this conjecture fails for the sparse subsequence of the form $N=2^{2^{w}-1}$, and the case $N=2^{2^{w}}$ was subsequently analyzed in [DGSxx]. Nevertheless, a complete characterization of the separation radius in the primary case $N=2^{m}$ remained open, and closing this gap is the aim of the present work.
@@
  • We provide an exact formula for the $\ell^{\infty}$-separation radius of $Q_{2^{m}}$ for all $m$.
@@
Otherwise, decomposing $m$ as $m=2^{v}+2^{w}+c$ with integers $v>w\geq 0$ and $0\leq c<2^{w}$ (note that the case $(w,c)=(v-1,2^{v-1}-1)$ is excluded since this falls into the first case as $m=2^{v+1}-1$), then
@@
$q_{\infty}(Q_{2^{m}})$ $\leq 2^{-3m/4-3/2}\qquad\text{for }m\neq 1,5,$ (3)
@@
Sobol’ sequences are most commonly and effectively used when the number of points is a power of two ($N=2^{m}$). Therefore, our results address the primary use case and are sufficient for the vast majority of practical applications. Despite this practical sufficiency, the exact values of the separation radius for arbitrary $N$ remain a natural mathematical question, which we investigate in Section 4.
Related work. In recent years, the quasi-uniformity of QMC point sets has been the subject of intensive investigation. There are two major classes of QMC constructions: digital nets and sequences [DP10, N92], and lattice point sets (including their infinite analogue, Kronecker sequences) [DKP22, SJ94].
For lattice point sets, quasi-uniformity has been actively studied. In one dimension, the separation radius of Kronecker sequences is completely characterized in [G24b]. More general results in higher dimensions are provided in [DGLPSxx], including bounded mesh ratios for two-dimensional Fibonacci lattices, existence results for $d$-dimensional lattice rules, and explicit constructions for $d$-dimensional Kronecker sequences.
Turning to digital nets and sequences, the one-dimensional van der Corput sequence in base $b$ is known to be quasi-uniform since its first $b^{m}$ points are $\{i/b^{m}\mid 0\leq i<b^{m}\}$. In higher dimensions, the covering radius, often referred to as dispersion, has been extensively studied [N92, Chapter 6]. In particular, for $(t,d)$-sequences in base bb—well-known examples include the Sobol’, Faure, and Niederreiter sequences—, the covering radius is known to attain the optimal order $\Theta(N^{-1/d})$ for any dimension $d$. Thus, the problem of establishing quasi-uniformity reduces to verifying whether the separation radius also scales as $N^{-1/d}$.
For $d=2$, as stated, the Sobol’ sequence is not quasi-uniform [G24a]. The separation radius of several two-dimensional digital nets was studied in [GHSK08, GK09]. Numerical experiments therein suggest that the Larcher–Pillichshammer nets [LP01] are quasi-uniform. This was theoretically proved by Dick, Goda, and Suzuki [DGSxx], who introduced an algebraic criterion for well-separated digital nets. To our knowledge, this remains the only explicit construction of low-discrepancy and quasi-uniform digital nets for $d\geq 2$. The paper [DGSxx] also shows the non-optimality of the separation radius for some two-dimensional digital nets and Fibonacci polynomial lattices, as well as for $b$-dimensional Faure sequences in prime base $b$.
From a different perspective, Pronzato and Zhigljavsky [PZ23] constructed quasi-uniform infinite sequences via a greedy packing algorithm, ensuring a mesh ratio of at most $2$ for the first $N$ points, $N\geq 2$. However, these sequences do not necessarily maintain low discrepancy in dimensions $d\geq 2$.
Organization. Preliminaries and notation are collected in Section 2. Section 3 provides the necessary lemmas, the proof of Theorem 1.1, and the derivation of Corollaries 1.2 and 1.3. Results for arbitrary NN are given in Section 4.
@@ 2 Preliminaries
For a vector $\vec{z}=(z_{1},\dots,z_{m})^{\top}\in\mathbb{F}_{2}^{m}$, we denote by $\vec{z}[i]\in\{0,1\}$ its $i$th component $z_{i}$, and for $i<j$ we define the slice $\vec{z}[i\!:\!j]=(z_{i},\dots,z_{j})^{\top}$. Finally, we set
$\phi(\vec{z})=\vec{z}[1]2^{-1}+\cdots+\vec{z}[m]2^{-m}.$
@@
Two-dimensional Sobol’ sequence. The two-dimensional Sobol’ sequence is defined as follows. For the definition in general dimension $d$, we refer the reader to [DP10, Chapter 8].
@@
The two-dimensional Sobol’ sequence $(\boldsymbol{x}_{n})_{n\in\mathbb{N}}$ is a sequence of points in $0,1)^{2}$, where the $n$th point is given by
$\boldsymbol{x}_{n}:=(\phi(\vec{n}),\phi(P_{m}\vec{n})),$
where $m\in\mathbb{N}$ is chosen such that $n<2^{m}$. This definition does not depend on the choice of $m$: increasing $m$ simply pads $\vec{n}$ with leading zeros, and since $P_{m}$ is upper triangular, $\phi(\vec{n})$ and $\phi(P_{m}\vec{n})$ remain unchanged.
Note that some implementations generate Sobol’ sequences in Gray code order rather than in the radical inverse order used here. This difference affects only the ordering of the points; for $N=2^{m}$, the resulting point sets are identical.
This construction is generalized to the notions of digital nets and digital sequences. The uniformity of these point sets is usually measured by a quantity called the $t$-value (see [N92, Chapter 4] or [DP10, Chapter 4]). It is known that the $t$-value of the two-dimensional Sobol’ sequence attains the best possible value, namely zero [DP10, Section 8.1]. This discussion can be formulated rigorously in the following proposition.
@@ 3.1 Lemmas
Using this fact, Items (vi)(viii) also follow from Lucas’s theorem. ∎
@@ 3.2 Proof of the main theorem
  • The case $m=2^{v}-1$ is essentially treated in [G24a]. As shown in the proof of [G24a, Theorem 2.2], we have
    $\|\boldsymbol{x}_{1}-\boldsymbol{x}_{2^{m}-2}\|_{\infty}=2^{-m}.$
@@
  • The case $m=2^{v}$ is essentially treated in [DGSxx]. From the proof of [DGSxx, Theorem 4.2] with $L$ being the identity matrix, we deduce $\boldsymbol{x}_{2^{m-1}+1}=(1/2+1/2^{m},1/2-1/2^{m})$, which implies $\|\boldsymbol{x}_{1}-\boldsymbol{x}_{2^{m-1}+1}\|_{\infty}=2^{-m}$. Hence, in the same manner as the previous case, we obtain
@@ 3.2.1 Proof of the upper bound
Let $p=2^{V+W-1}+2^{V-1}$ and $q=2^{V+W}-2^{V}$. To prove $q_{\infty}(Q_{2^{m}})\leq 2^{-V-W}$, it suffices to show that $\|\boldsymbol{x}_{p}-\boldsymbol{x}_{q}\|_{\infty}=2^{-V-W+1}$.
First, we compute $\boldsymbol{x}_{p}=(x_{p,1},x_{p,2})$. Since $\vec{p}[i]=1$ if and only if $i=V$ or $i=V+W$, it follows from Lemma 3.1 (i) and (ii) that
$x_{p,1}$ $=\phi(\vec{p})=2^{-V}+2^{-V-W},$
$x_{p,2}$ $=\phi(P_{m}\vec{p})=\sum_{i=1}^{V+W}(P[i][V]\oplus P[i][V+W])2^{-i}=\sum_{i=W+1}^{V+W}2^{-i}=2^{-W}-2^{-V-W}.$
Next, we compute $\boldsymbol{x}_{q}=(x_{q,1},x_{q,2})$. Since $\vec{q}[i]=1$ for $V+1\leq i\leq V+W$, we have
$P_{m}\vec{q}$ $=P_{m}(\mathrm{concat}(\boldsymbol{1}_{V},\boldsymbol{0}_{m-V})\oplus\mathrm{concat}(\boldsymbol{1}_{V+W},\boldsymbol{0}_{m-V-W}))$
$=\mathrm{concat}(P_{V}\boldsymbol{1}_{V},\boldsymbol{0}_{m-V})\oplus\mathrm{concat}(P_{V+W}\boldsymbol{1}_{V+W},\boldsymbol{0}_{m-V-W}),$
@@ 3.2.2 Proof of the lower bound
Then, Corollary 3.3 (i) with $\ell=V+1$ implies
@@
In this case, Corollary 3.3 (ii) with $k=V$ and $\ell=V+W$, together with the assumption $\phi(\vec{p})<\phi(\vec{q})$, implies
@@
In this case, using (12), (8), Lemma 3.1 (v), and Lemma 3.1 (viii), we have
$\boldsymbol{1}$ $=(P\vec{\Delta})[V+1\!:\!V+W-1]$
$=P[V+1\!:\!V+W-1][V+1\!:\!V+W-1]\cdot\vec{\Delta}[V+1\!:\!V+W-1]$
$\qquad\oplus P[V+1\!:\!V+W-1][V+W\!:\!V+W]\cdot\vec{\Delta}[V+W]$
$\qquad\oplus P[V+1\!:\!V+W-1][V+W+1\!:\!m]\cdot\vec{\Delta}[V+W+1\!:\!m]$
@@
$=\boldsymbol{1}\oplus P^{\prime}\cdot\vec{\Delta}[V+W+1\!:\!m].$
where $P^{\prime}:=P[V+1\!:\!V+W-1][V+W+1\!:\!m]$. This gives
@@
Finally, (5) follows immediately from (4) together with the general bound $h_{\infty}(Q_{N})\geq 1/(2\sqrt{N})$ as given in [DGSxx, Remark 2.4]. ∎
4 Separation radius for general NN
Building on our previous result for the exact separation radius $q_{\infty}(Q_{2^{m}})$ for a power-of-two number of points, this section establishes an upper bound for $q_{\infty}(Q_{N})$ for an arbitrary number of points $N$. It is important to note that the analysis herein applies exclusively to the Sobol’ sequence in radical inverse order (Definition 2.1). To motivate our theoretical result, we first conducted a numerical experiment to identify the pairs of indices $(p,q)$ that successively update the minimum distance $\|\boldsymbol{x}_{p}-\boldsymbol{x}_{q}\|_{\infty}$ for $q\leq 2^{15}$. Note that, since the Sobol’ sequence starts from index 0, the point set corresponding to $(p,q)$ consists of $N=q+1$ points. We recall that the separation radius is the half of the minimum distance. The results of this search, summarized in Table 1, motivate the following theorem, which provides an analytical result consistent with our numerical observations.
Table 1: Indices $(p,q)$ for which the distance $\|\boldsymbol{x}_{p}-\boldsymbol{x}_{q}\|_{\infty}$ successively achieves the minimum distance for the Sobol’ sequence, shown for $q\leq 2^{15}$. The fourth row indicates the corresponding case in Theorem 4.1 that describes the pair. The last two rows show the corresponding number of points $N=q+1$ and the separation radius $q_{\infty}(Q_{N})=\|\boldsymbol{x}_{p}-\boldsymbol{x}_{q}\|_{\infty}/2$, respectively.
pp 0 11 22 11 $24$ $40$ 88 11 $504$ $640$ $2176$ $128$
$q$ 11 22 44 99 $36$ $48$ $112$ $129$ $516$ $768$ $3840$ $32512$
$\|\boldsymbol{x}_{p}-\boldsymbol{x}_{q}\|_{\infty}$ $^{-1} 22$$^{-2}$ $2^{-3}$ $2^{-4}$ $3\cdot 2^{-6}$ $2^{-5}$ $2^{-7}$ $2^{-8}$ $3\cdot 2^{-10}$ $2^{-9}$ $2^{-11}$ $2^{-15}$
Thm. 4.1 (ii) (i) (iii) (iv) (ii) (i) (iii) (iv) (iv) (ii)
NN 22 33 55 $10$ $37$ $49$ $113$ $130$ $517$ $769$ $3841$ $32513$
$q_{\infty}(Q_{N})$ $2^{-2}$ $2^{-3}$ $2^{-4}$ $2^{-5}$ $3\cdot 2^{-7}$ $2^{-6}$ $2^{-8}$ $2^{-9}$ $3\cdot 2^{-11}$ $2^{-10}$ $2^{-12}$ $2^{-16}$
Theorem 4.1.
Let $Q=\{\boldsymbol{x}_{0},\boldsymbol{x}_{1},\dots\}$ be the two-dimensional Sobol’ sequence and $Q_{N}$ be its first $N$ points. Let $v>w\geq 1$ be integers. Then we have the following.
  1. (i)
    Case where $m=2^{v}$: Let $p=1$, $q=2^{2^{v}-1}+1$, and $N=q+1$. Then
    $q_{\infty}(Q_{N})\leq\dfrac{1}{2}\|\boldsymbol{x}_{p}-\boldsymbol{x}_{q}\|_{\infty}=2^{-2^{v}-1}$
  2. (ii)
    Case where $m=2^{v}-1$: Let $p=2^{2^{v-1}-1}$, $q=2^{2^{v}-1}-2^{2^{v-1}}$, and $N=q+1$. Then
    $q_{\infty}(Q_{N})\leq\dfrac{1}{2}\|\boldsymbol{x}_{p}-\boldsymbol{x}_{q}\|_{\infty}=2^{-2^{v}}.$
  3. (iii)
    Case where $m=2^{v}+2$: Let $p=2^{2^{v}+1}-8$, $q=2^{2^{v}+1}+4$, and $N=q+1$. Then
    $q_{\infty}(Q_{N})\leq\dfrac{1}{2}\|\boldsymbol{x}_{p}-\boldsymbol{x}_{q}\|_{\infty}=3\cdot 2^{-2^{v}-3}$
  4. (iv)
    Case where $m=2^{v}+2^{w}$: Let $p=2^{2^{v}+2^{w}-1}+2^{2^{v}-1}$, $q=2^{2^{v}+2^{w}}-2^{2^{v}}$, and $N=q+1$. Then
    $q_{\infty}(Q_{N})\leq\dfrac{1}{2}\|\boldsymbol{x}_{p}-\boldsymbol{x}_{q}\|_{\infty}=2^{-2^{v}-2^{w}}.$
Proof.
In this proof, we use the notation $V=2^{v}$ and $W=2^{w}$ as in Section 3.
(i): This is shown in the second case of Section 3.2.
(ii): Let $V^{\prime}=2^{v-1}$. First, we compute $\boldsymbol{x}_{p}=(x_{p,1},x_{p,2})$. Since $\vec{p}[i]=1$ if and only if $i=V^{\prime}$ , it follows from Lemma 3.1 (i) that
$x_{p,1}$ $=\phi(\vec{p})=2^{-V^{\prime}},$
$x_{p,2}$ $=\phi(P_{V-1}\vec{p})=\sum_{i=1}^{V^{\prime}}P[i][V^{\prime}]2^{-i}=\sum_{i=1}^{V^{\prime}}2^{-i}=1-2^{-V^{\prime}}.$
Next, we compute $\boldsymbol{x}_{q}=(x_{q,1},x_{q,2})$. In what follows, “concat” denotes vertical concatenation of column vectors. Since $\vec{q}[i]=1$ for $V^{\prime}+1\leq i\leq V-1$, we have
$P_{V-1}\vec{q}$ $=P_{V-1}(\boldsymbol{1}_{V-1}\oplus\mathrm{concat}(\boldsymbol{1}_{V^{\prime}},\boldsymbol{0}_{V-V^{\prime}-1}))$
$=P_{V-1}\boldsymbol{1}_{V-1}\oplus\mathrm{concat}(P_{V^{\prime}}\boldsymbol{1}_{V^{\prime}},\boldsymbol{0}_{V-V^{\prime}-1}))$
[…]
$=\mathrm{concat}(\boldsymbol{1}_{V^{\prime}-1},\boldsymbol{0}_{1},\boldsymbol{1}_{V-V^{\prime}-1}),$
where the third equality follows from Lemma 3.1 (vi) and (viii). Thus we obtain
$x_{q,1}$ $=\phi(\vec{q})=\sum_{i=V^{\prime}+1}^{V-1}2^{-i}=2^{-V^{\prime}}-2^{-V+1}$
$x_{q,2}$ $=\phi(P_{V-1}\vec{q})=\sum_{i=1}^{V-1}2^{-i}-2^{-V^{\prime}}=1-2^{-V^{\prime}}-2^{-V+1}.$
Hence $\|\boldsymbol{x}_{p}-\boldsymbol{x}_{q}\|=2^{-V+1}$.
(iii): First, we compute $\boldsymbol{x}_{p}=(x_{p,1},x_{p,2})$. We have $\vec{p}[i]=1$ for $4\leq i\leq V+1$. From Lemma 3.1 (vii) with $(V,W)=(2,1)$ and $(V,1)$, we have
$(P_{3}\boldsymbol{1}_{3})[i]$ $=1\iff i=1,2,3$
$(P_{V+1}\boldsymbol{1}_{V+1})[i]$ $=1\iff i=1,V,V+1.$
Since we have assumed that $v\geq 2$ and thus $V\geq 4$, we have
$(P_{V+2}\vec{p})[i]=1\iff i=2,3,V,V+1,$
and thus
$x_{p,1}$ $=\phi(\vec{p})=\sum_{i=4}^{V+1}2^{-i}=2^{-3}-2^{-V-1},$
$x_{p,2}$ $=\phi(P_{V+2}\vec{p})=2^{-2}+2^{-3}+2^{-V}+2^{-V-1}.$
Next, we compute $\boldsymbol{x}_{q}=(x_{q,1},x_{q,2})$. Since $\vec{q}[i]=1$ if and only if $i=3$ or $i=V+2$, it follows from Lemma 3.1 (ii) with $(V,W)=(2,1)$ and $(V,2)$ that
$x_{q,1}$ $=\phi(\vec{q})=2^{-3}+2^{-V-2}$
$x_{q,2}$ $=\phi(P_{V+2}\vec{q})=\sum_{i=1}^{V+2}(P[i][3]\oplus P[i][V+2])2^{-i}=2^{-2}+2^{-3}+2^{-V-1}+2^{-V-2}.$
Hence $\|\boldsymbol{x}_{p}-\boldsymbol{x}_{q}\|=3\cdot 2^{-V-2}$.
(iv): This is shown in Section 3.2.1. ∎
Establishing a lower bound on the separation radius or minimum distance for general $N$ would necessitate a significantly more complex argument than the one used for the proof in Section 3.2.2. We therefore propose the conjecture that, for $q\geq 4$, a pair $(\boldsymbol{x}_{p},\boldsymbol{x}_{q})$ updates the minimum distance if and only if it is one of the candidates specified in Theorem 4.1.
5 Conclusion
We established a complete characterization of the separation radius for the case $N=2^{m}$ and investigated its numerical behavior for general $N$. Our analysis reveals that the two-dimensional Sobol’ sequence has a suboptimal mesh ratio that grows at least as $N^{1/4}$. While digital scrambling may alleviate the poor growth of the separation radius, examining this possibility is left for future work.
In sharp contrast to the two-dimensional case (Theorem 1.1), Sobol’ points in dimensions $d\geq 3$ did not exhibit a clear systematic pattern in their separation radius. For $d=3$, in particular, the observed rate lies between $N^{-1/3}$ and $N^{-1/2}$ up to $N\leq 2^{15}$. Understanding the behavior in higher dimensions and establishing dimension-dependent bounds remain important directions for future research.
Acknowledgments
The author acknowledge the use of OpenAI’s ChatGPT 5 and Google’s Gemini 2.5 for editing and language polishing of the manuscript.